# Thermodynamics of EMD-AdS black hole with deformed horizon

### v1

@@ Abstract
[a]organization=Department of Physics,addressline=South China Normal University, city=Guangzhou, postcode=510006, country=China
@@
[b]organization=Guangdong Provincial Key Laboratory of Quantum Engineering and Quantum Materials, Guangdong-Hong Kong Joint Laboratory of Quantum Matter,addressline=South China Normal University, city=Guangzhou, postcode=510006, country=China
@@
[c]organization=Department of Astronomy, School of Physics and Materials Science,addressline=Guangzhou University, city=Guangzhou, postcode=510006, country=China
@@ 1 Introduction
One class of AdS black holes are associated with EMD systems. These systems, which couple a Maxwell field and a dilaton field, have received widespread attention due to their ability to holographically describe systems with finite charge density or condensed matter [6, 7, 8, 9, 10, 11, 12]. For instance, the holographic description of these black holes has been used to explore the properties of strongly coupled field theories at finite temperature and density.
@@
On the other hand, AdS background permits a black hole with topologically non-spherical horizon [22, 23, 24, 25]. Furthermore, the maximal symmetry can be broken by introducing holographic axions [26, 27, 28, 29, 30, 31]. Some of these solutions constructed by the method proposed in Refs.[32, 33] allow the Ricci scalar of the angular space $\hat{R}$ to be non-constant, instead depending on angular coordinate. The thermodynamics of topological black holes has also been extensively studied [34, 35, 26]. We will take one of these solutions as an example to demonstrate that the angular part of such deformed solutions can be combined with the radial part of EMD black holes through a warped product, forming a new metric that satisfies the Einstein equations. The interesting aspect of these new solutions is that the deformation does not alter the action and thermodynamic quantities of the EMD black holes, but only changes the curvature of the angular space. Referring to the unified first law [36, 37, 38], we introduce a quasi-local method to study the thermodynamics of the deformed solutions. This approach is similar to the one used in the context of perfect fluid dark matter black holes in a phantom background, where the MS mass [39, 40] is identified as the total energy of the spacetime within a certain radius, and it helps in formulating the corresponding thermodynamic first law.
@@ 2 Einstein-Maxwell-Dilaton Black Hole with Axions
$\begin{split}\mathcal{L}=&\frac{R}{16\pi G}-\frac{1}{4}e^{-\alpha\phi}F^{\mu\nu}F_{\mu\nu}\\ &-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\mathcal{V}(\phi)-\frac{1}{2}\sum_{i=1}^{2}\nabla_{\mu}\psi^{i}\nabla^{\mu}\psi^{i}\,,\end{split}$ (2)
@@
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}{Rg_{\mu\nu}}=8\pi\left({T^{(em)}}_{\mu\nu}+{T^{(\phi)}}_{\mu\nu}+{T^{(\psi^{i})}}_{\mu\nu}\right)\,,$ (4)
$\nabla^{\mu}F_{\mu\nu}=0\,,$ (5)
$\nabla^{\mu}\nabla_{\mu}\phi+\frac{\partial{\cal L}}{\partial\phi}=0\,,$ (6)
@@
${T^{(em)}}_{\mu\nu}=\frac{1}{16\pi}{e^{-\alpha\phi(r)}\left(F_{\mu\sigma}{F_{\nu}}^{\sigma}-\frac{1}{4}F_{\rho\sigma}F^{\rho\sigma}g_{\mu\nu}\right)}\,,$ (8)
${T^{(\phi)}}_{\mu\nu}=\frac{1}{16\pi}\left({\partial_{\mu}}{\phi}{\partial_{\nu}}{\phi}-g_{\mu\nu}\left(\frac{1}{2}{\partial_{\alpha}\phi}\partial^{\alpha}\phi+\mathcal{V}(\phi)\right)\right)\,,$ (9)
${T^{({\psi}^{i})}}_{\mu\nu}=\frac{1}{16\pi}({\partial_{\mu}}{{\psi}^{i}}{\partial_{\nu}}{{\psi}^{i}}-\frac{1}{2}g_{\mu\nu}{\partial_{\alpha}{\psi}^{i}}\partial^{\alpha}{\psi}^{i})\,.$ (10)
@@
$\begin{split}V(\phi)+&\frac{1}{2}e^{-\alpha\phi}{A_{t}}^{\prime 2}+\frac{f^{\prime}U^{\prime}+fU^{\prime\prime}-2\text{k}}{U}\\ &+\frac{1}{2U}(-\frac{2}{\rho}\frac{de}{d{\rho}}+(1-k\rho^{2}-e)(\frac{d{\psi}^{1}}{d{\rho}})^{2}+\frac{1}{{\rho^{2}}}(\frac{d{\psi}^{2}}{d{\varphi}})^{2})=0\,,\end{split}$ (14)
$e^{-\alpha\phi}{A_{t}^{\prime}}^{2}-f^{\prime\prime}+\frac{fU^{\prime\prime}-2k}{U}=0\,,$ (15)
@@
$\begin{cases}-\frac{2}{\rho}\frac{de}{d{\rho}}+(1-k\rho^{2}-e)(\frac{d{\psi}^{1}}{d{\rho}})^{2}+\frac{1}{{\rho^{2}}}(\frac{d{\psi}^{2}}{d{\varphi}})^{2}=-\eta\\ \mathcal{V}(\phi)U+\frac{1}{2}e^{-\alpha\phi}{A_{t}}^{\prime 2}U+f^{\prime}U^{\prime}+fU^{\prime\prime}-2\text{k}=2\eta\end{cases}\,.$ (20)
@@
$\begin{cases}e(\rho)={\xi}^{2}\log(\rho)-\eta\rho^{2}+Const.\\ \frac{d{\psi}^{1}}{d{\rho}}=\frac{\xi}{\rho\sqrt{1-k{\rho}^{2}-e}}\end{cases}\,.$ (21)
We can consider the constant in $e(\rho)$ as $e(\rho)={\xi}^{2}\ln{(\zeta\rho)}-\eta\rho^{2}$. These solutions can satisfy equation (19) automatically. The analytic solutions of ${f(r),\phi(r),U(r),A_{t}(r)}$ are:
$f=\left(\text{k}+\eta-\frac{c}{r}\right)\left(1-\frac{b}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{r^{2}}{L^{2}}\left(1-\frac{b}{r}\right)^{\frac{2\alpha^{2}}{\alpha^{2}+1}}\,,$ (22)
@@
$\begin{split}ds^{2}=-(k-&\frac{2M}{r}+\frac{r^{2}}{L^{2}})dt^{2}+(k-\frac{2M}{r}+\frac{r^{2}}{L^{2}})^{-1}dr^{2}\\ &+r^{2}(\frac{d\rho^{2}}{1-k\rho^{2}-e(\rho)}+\rho^{2}d\varphi^{2})\,.\end{split}$ (30)
@@ 4 Thermodynamics and Unified First Law
$\begin{split}M_{MS}(u,x)=&\frac{(D-2)\Omega}{16\pi G}\tilde{r}^{D-3}(\frac{\hat{R}}{(D-2)(D-3)}-\bar{g}^{\tilde{r}\tilde{r}})\\ &-\frac{\Lambda}{8\pi G}\frac{\Omega\,\tilde{r}^{D-1}}{D-1}\,,\end{split}$ (43)
@@
$\begin{split}M_{MS}(\tilde{r},\rho)=\frac{\Omega}{8\pi G}[\tilde{r}(\frac{\hat{R}}{2}-f(\tilde{r})(\frac{d\tilde{r}}{dr})^{2})+{\tilde{r}}^{3}]\,,\end{split}$ (47)
@@
We propose to incorporate an axion charge term into the first law of thermodynamics while ensuring the validity of the equation. The first method is to apply quasi-local viewpoint and unified first law to reengineer the equation $dM_{0}-TdS=0$. We need to define the work term $W$ first in our ansatz
$W=-\frac{1}{2}T_{ab}g^{ab}\,.$ (48)
The equation of first law becomes
$dM_{MS}-WdV=TdS-\frac{\mathcal{V}(\phi)}{8\pi G}dV\,,$ (49)
which is based on the unified first law and relation between surface gravities and temperature. It implies that
$dM_{MS}-W_{H}dV_{H}+\frac{\mathcal{V}(\phi)}{8\pi G}dV_{H}=dM\,,$ (50)
where we use subscript $H$ to denote quasi-local physical quantities at the horizon. This equation implies that we can study the connection between the quasi-local viewpoint with the global viewpoint.
@@
$M_{MS}(\tilde{r_{h}},\rho)=\frac{\Omega}{8\pi G}[k\tilde{r_{h}}+\tilde{r_{h}}^{3}+\frac{\xi^{2}}{2\rho^{2}}\tilde{r_{h}}]\,.$ (51)
@@
In the situation without angular deformation, the MS mass will be
$M_{0}({r_{h}})=\frac{\Omega kr_{h}}{8\pi G}\left(1-\left(1-\frac{b}{{r_{h}}}\right)^{\frac{1-\alpha^{2}}{\alpha^{2}+1}}\right)\left(1-\frac{b}{{r_{h}}}\right)^{\frac{\alpha^{2}}{\alpha^{2}+1}}\,,$ (53)
which is a function of ${r_{h}}$ and k, similar to the temperature. When $r\rightarrow\infty$, it will be
$M_{0,r\rightarrow\infty}=\frac{\Omega k}{8\pi G}(\frac{1-\alpha^{2}}{1+\alpha^{2}}b)\,,$ (54)
which returns to equation (40) in the hyperbolic situation $k=-1$.
When there are axions in EMD-ADS spacetime, we see the first law as [42]
$dM+\varpi d\Pi+PdV=TdS\,,$ (55)
@@
$\begin{split}P_{\xi}&=(\frac{\partial M_{\xi}}{\partial V})_{\varpi}\\ &=\frac{3a^{2}}{16\pi GU(r)\theta^{2}}\,.\end{split}$ (57)
@@
$\begin{split}\Pi&=\frac{1}{L}\sqrt{\frac{1}{D-2}\sum(\partial\psi^{i})^{2}}\\ &=\frac{a}{2\theta U^{\frac{1}{2}}}\,.\end{split}$ (58)
@@
$\varpi=-\frac{a(-k)^{\frac{2\alpha^{2}}{3\alpha^{2}-1}}r^{\frac{2\left(\alpha^{2}-1\right)}{3\alpha^{2}-1}}}{8\pi G\theta}\Omega\,.$ (59)

### v2

@@ Abstract
[1]organization=Department of Physics,addressline=South China Normal University, city=Guangzhou, postcode=510006, country=China
@@
[2]organization=Guangdong Provincial Key Laboratory of Quantum Engineering and Quantum Materials, Guangdong-Hong Kong Joint Laboratory of Quantum Matter,addressline=South China Normal University, city=Guangzhou, postcode=510006, country=China
@@
[3]organization=Department of Astronomy, School of Physics and Materials Science,addressline=Guangzhou University, city=Guangzhou, postcode=510006, country=China
@@ 1 Introduction
One class of AdS black holes are associated with EMD systems. These systems, which couple a Maxwell field and a dilaton field, have received widespread attention due to their ability to holographically describe systems with finite charge density or condensed matter [6, 7, 8, 9, 10, 11, 12, 13, 14]. For instance, the holographic description of these black holes has been used to explore the properties of strongly coupled field theories at finite temperature and density.
@@
On the other hand, AdS background permits a black hole with topologically non-spherical horizon [24, 25, 26, 27]. Furthermore, the maximal symmetry can be broken by introducing holographic axions [28, 29, 30, 31, 32, 33]. Some of these solutions constructed by the method proposed in Refs.[34, 35] allow the Ricci scalar of the angular space $\hat{R}$ to be non-constant, instead depending on angular coordinate.
The thermodynamics of topological black holes has also been extensively studied [13, 36, 37, 28, 38, 39, 40]. We will take one of these solutions as an example to demonstrate that the angular part of such deformed solutions can be combined with the radial part of EMD black holes through a warped product, forming a new metric that satisfies the Einstein equations. The interesting aspect of these new solutions is that the deformation does not alter the action and thermodynamic quantities of the EMD black holes, but only changes the curvature of the angular space. Referring to the unified first law [41, 42, 43], we introduce a quasi-local method to study the thermodynamics of the deformed solutions. This approach is similar to the one used in the context of perfect fluid dark matter black holes in a phantom background, where the MS mass [44, 45] is identified as the total energy of the spacetime within a certain radius, and it helps in formulating the corresponding thermodynamic first law.
@@ 2 Einstein-Maxwell-Dilaton Black Hole with Axions
$\begin{split}\mathcal{L}=&\frac{R}{16\pi G}-\frac{1}{16}e^{-\alpha\phi}F^{\mu\nu}F_{\mu\nu}\\ &-\frac{1}{2G}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{G}\mathcal{V}(\phi)-\frac{1}{2}\sum_{i=1}^{2}\nabla_{\mu}\psi^{i}\nabla^{\mu}\psi^{i}\,,\end{split}$ (2)
@@
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}{Rg_{\mu\nu}}=8\pi G\left({T^{(em)}}_{\mu\nu}+{T^{(\phi)}}_{\mu\nu}+{T^{(\psi^{i})}}_{\mu\nu}\right)\,,$ (4)
$\nabla^{\mu}(e^{-\alpha\phi}F_{\mu\nu})=0\,,$ (5)
$\nabla^{\mu}\nabla_{\mu}\phi=\frac{\partial{\mathcal{V}}}{\partial\phi}\,,$ (6)
@@
${T^{(em)}}_{\mu\nu}={e^{-\alpha\phi}\left(F_{\mu\sigma}{F_{\nu}}^{\sigma}-\frac{1}{4}F_{\rho\sigma}F^{\rho\sigma}g_{\mu\nu}\right)}\,,$ (8)
${T^{(\phi)}}_{\mu\nu}=\frac{1}{G}[{\partial_{\mu}}{\phi}{\partial_{\nu}}{\phi}-g_{\mu\nu}\left(\frac{1}{2}{\partial_{\alpha}\phi}\partial^{\alpha}\phi+\mathcal{V}(\phi)\right)]\,,$ (9)
${T^{({\psi}^{i})}}_{\mu\nu}=({\partial_{\mu}}{{\psi}^{i}}{\partial_{\nu}}{{\psi}^{i}}-\frac{1}{2}g_{\mu\nu}{\partial_{\alpha}{\psi}^{i}}\partial^{\alpha}{\psi}^{i})\,.$ (10)
@@
$\begin{split}V(\phi)&+8\pi Ge^{-\alpha\phi}{A_{t}}^{\prime 2}+\frac{f^{\prime}U^{\prime}+fU^{\prime\prime}-2\text{k}}{U}\\ &+\frac{1}{2U}(-\frac{2}{\rho}\frac{de}{d{\rho}}+16\pi G(1-k\rho^{2}-e)(\frac{d{\psi}^{1}}{d{\rho}})^{2}+\frac{16\pi G}{{\rho^{2}}}(\frac{d{\psi}^{2}}{d{\varphi}})^{2})=0\,,\end{split}$ (14)
$16\pi Ge^{-\alpha\phi}{A_{t}^{\prime}}^{2}-f^{\prime\prime}+\frac{fU^{\prime\prime}-2k}{U}=0\,,$ (15)
@@
$\begin{cases}-\frac{2}{\rho}\frac{de}{d{\rho}}+16\pi G(1-k\rho^{2}-e)(\frac{d{\psi}^{1}}{d{\rho}})^{2}+\frac{16\pi G}{{\rho^{2}}}(\frac{d{\psi}^{2}}{d{\varphi}})^{2}=-\eta\\ \mathcal{V}(\phi)U+8\pi Ge^{-\alpha\phi}{A_{t}}^{\prime 2}U+f^{\prime}U^{\prime}+fU^{\prime\prime}-2\text{k}=2\eta\end{cases}\,.$ (20)
@@
$\begin{cases}e(\rho)=16\pi G{\xi}^{2}\log(\rho)-\eta\rho^{2}+Const.\\ \frac{d{\psi}^{1}}{d{\rho}}=\frac{\xi}{\rho\sqrt{1-k{\rho}^{2}-e}}\end{cases}\,.$ (21)
We can consider the constant in $e(\rho)$ as $e(\rho)=16\pi G{\xi}^{2}\log{(\zeta\rho)}-\eta\rho^{2}$. These solutions can satisfy equation (19) automatically. The analytic solutions of ${f(r),\phi(r),U(r),A_{t}(r)}$ are:
$f=\left(\text{k}+\eta-\frac{16\pi Gc}{r}\right)\left(1-\frac{b}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{r^{2}}{L^{2}}\left(1-\frac{b}{r}\right)^{\frac{2\alpha^{2}}{\alpha^{2}+1}}\,,$ (22)
@@
$\begin{split}ds^{2}=-(k-&\frac{16\pi Gc}{r}+\frac{r^{2}}{L^{2}})dt^{2}+(k-\frac{16\pi Gc}{r}+\frac{r^{2}}{L^{2}})^{-1}dr^{2}\\ &+r^{2}(\frac{d\rho^{2}}{1-k\rho^{2}-e(\rho)}+\rho^{2}d\varphi^{2})\,.\end{split}$ (30)
@@ 4 Thermodynamics and Unified First Law
$M_{MS}(u,x)=\frac{(D-2)\Omega}{16\pi G}\tilde{r}^{D-3}(\frac{\hat{R}}{(D-2)(D-3)}-\bar{g}^{\tilde{r}\tilde{r}})\,,$ (43)
@@
$\begin{split}M_{MS}(\tilde{r},\rho)=\frac{\Omega}{8\pi G}\tilde{r}\bigg{(}\frac{\hat{R}}{2}-f(\tilde{r})\tilde{r}^{\prime 2}\bigg{)}\,,\end{split}$ (47)
@@
We propose to incorporate an axion charge term into the first law of thermodynamics while ensuring the validity of the equation. The first method is to apply quasi-local viewpoint and unified first law to reengineer the equation $dM-TdS=0$. We need to define the energy supply vector $\Psi^{a}$ and the work term $W$ first in our ansatz
$\begin{split}\Psi^{a}&=(T^{ab}-\frac{1}{2}T_{cd}\bar{g}^{cd}\bar{g}^{ab})\bar{\nabla}_{b}\tilde{r}=-\frac{1}{8\pi G}f(r)\frac{\tilde{r}^{\prime}\tilde{r}^{\prime\prime}}{\tilde{r}}\bar{\nabla}_{a}\tilde{r},\\ W&=-\frac{1}{2}T_{ab}\bar{g}^{ab}=-\frac{1}{8\pi G\tilde{r}}\left(\bar{\nabla}^{2}\tilde{r}-\frac{1}{\tilde{r}}(\frac{\hat{R}}{2}-\bar{g}^{\tilde{r}\tilde{r}})\right)\,,\end{split}$ (48)
where $T_{\mu\nu}={T^{(em)}}_{\mu\nu}+{T^{(\phi)}}_{\mu\nu}+{T^{(\psi^{i})}}_{\mu\nu}$ is the total energy-momentum tensor. The equation of unified first law becomes
$(dM_{MS})_{a}=W(dV)_{a}+A\Psi_{a}\,,$ (49)
where the ”volume” $V=\Omega r^{D-1}/(D-1)$ and area $A=\Omega r^{D-2}$. If let them take values on the horizon, the unified first law implies the first law of black hole thermodynamics in terms of such quasi-local parameters. The unified first law yields
$2\pi\frac{\delta M_{MS}-W_{H}\delta V_{H}}{(\kappa_{geo})_{H}}=\frac{\delta A_{H}}{4G}=\delta S\,,$ (50)
where we use subscript $H$ to denote quasi-local physical quantities at the horizon and the surface gravity is $\kappa_{geo}\equiv\frac{1}{2}\bar{\nabla}^{2}\tilde{r}$. This equation implies that we can study the connection between the quasi-local viewpoint with the global viewpoint.
@@
$M_{MS}(\tilde{r}_{h},\rho)=\frac{\Omega\tilde{r}_{h}}{8\pi G}[k+\frac{8\pi G\xi^{2}}{\rho^{2}}]\,.$ (51)
@@
When there are axions in EMD-ADS spacetime, we see the first law as [47]:
$dM_{MS}+\varpi d\Pi+PdV=dM_{0}\,,$ (53)
@@
$P_{\xi}=(\frac{\partial M_{\xi}}{\partial V})_{\varpi}=\frac{3a^{2}}{U(r)\theta^{2}}\,.$ (55)
@@
$\Pi=\frac{1}{L}\sqrt{\frac{1}{D-2}\sum(\partial\psi^{i})^{2}}=\frac{a}{2\theta U^{\frac{1}{2}}}\,.$ (56)
@@
$\varpi=-\frac{a(-k)^{\frac{2\alpha^{2}}{3\alpha^{2}-1}}r^{\frac{2\left(\alpha^{2}-1\right)}{3\alpha^{2}-1}}}{\theta}\Omega\,.$ (57)